# Counting Matrices in $\mathrm{SL}_{3}(\mathbb{Z})$ with Fixed Completely Split Character Polynomial: Preliminary Upper Bounds

### v1

@@ 2.3 Counting
This naive bound is $O(H^{4})$; we refine it by observing we overcounted $b$: in fact $b$ itself must satisfy $|b|\leq H$, independent of $a,c,u$ (since $B_{13}$ contains $b$ with coefficient $1$ plus smaller terms). Thus we replace the preceding factor $H$ by $1$, obtaining
$N_{(x+1)^{2}(x-1)}(H)\ll\sum_{|a|,|c|\leq H}\left(\frac{H}{1+|a|+|c|}\right)^{3}.$ (6)
Now write $m=1+|a|+|c|$ and note that the number of pairs $(a,c)$ with given $m$ is $\ll m$ (since $|a|,|c|\leq m$ and $|a|+|c|=m-1$ gives at most $m$ solutions). Hence
$N_{(x+1)^{2}(x-1)}(H)\ll\sum_{m\leq 2H+2}m\left(\frac{H}{m}\right)^{3}=H^{3}\sum_{m\leq 2H+2}\frac{1}{m^{2}}\ll H^{3}.$ (7)

### v2

@@ 2.3 Counting
This naive bound is $O(H^{4})$; we refine it by observing we overcounted $b$: in fact $b$ itself must satisfy $|b|\leq H$, independent of $a,c,u$ (since $B_{13}$ contains $b$ with coefficient $1$ plus smaller terms).
Refining the $b$–count.
For fixed $(a,c,u_{21},u_{31},u_{32})$ we have
$B_{12}=a-b\,u_{32}.$
If $u_{32}=0$, then $B_{12}=a$ and any $|b|\leq H$ is admissible, giving a factor $\ll H$. If $u_{32}\neq 0$, the constraint $|a-bu_{32}|\leq H$ implies
$|b|\leq\frac{|a|+H}{|u_{32}|}\quad\Rightarrow\quad\#\{b\}\ \ll\ \frac{H}{|u_{32}|}.$
From Lemma 2.3 we already know $|u_{32}|\ll H/(1+|a|+|c|)$, hence
[…]
Thus the crude factor $H$ is replaced by
$\min\!\left\{H,\frac{H}{|u_{32}|}\right\}=O\!\left(\frac{H}{1+|a|+|c|}\right)$
whenever $u_{32}\neq 0$. Summing over $u_{21},u_{31},u_{32}$ (whose ranges are $\ll H/(1+|a|+|c|)$) and then over $a,c$ gives
$\sum_{|a|,|c|\leq H}\left(\frac{H}{1+|a|+|c|}\right)^{3}=O(H^{3}),$
so the total remains $O(H^{3})$.
Hence the crude factor $H$ for $b$ is replaced by $O\!\left(H/(1+|a|+|c|)\right)$. Therefore
$N_{(x+1)^{2}(x-1)}(H)\ll\sum_{|a|,|c|\leq H}\left(\frac{H}{1+|a|+|c|}\right)^{3}.$
Now write $m=1+|a|+|c|$; the number of pairs $(a,c)$ with given $m$ is $\ll m$, so
$N_{(x+1)^{2}(x-1)}(H)\ll\sum_{m\leq 2H+2}m\left(\frac{H}{m}\right)^{3}=H^{3}\sum_{m\leq 2H+2}\frac{1}{m^{2}}\ll H^{3}.$